OpenStudy (anonymous):

let f(x)= 2/1+x^2. The equation of the tangent line to the curve at the point (5,0.07692) can be written in the form y=mx+b where m is: ? and b is?

OpenStudy (anonymous):

m = 2x at x=5

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

y1 x1 are your points plug, rearrange, and determine your values

OpenStudy (turingtest):

point-slope form for algebra is this:\[y-y_1=m(x-x_1)\]now in calculus we learn that slope of the tangent at a point 'a' is the derivative at that point\[m=f'(a)\]so we need the coordinates of the point\[(x_1,y_1)\]and the value of the derivative \[f'(x_1)\]then we can use the modified point-slope form\[y-y_1=f'(x_1)(x-x_1)\]

OpenStudy (turingtest):

rearrange into slope intercept form...